OpenStudy (anonymous):

A dies is thrown six times. (a) what is the probability that the nth throw is n on each occasion? (b) what is the probability that the nth throw is n on exactly five occasions?

OpenStudy (akshay_budhkar):

the first answer is (1/6)^6...

OpenStudy (akshay_budhkar):

i am not quite sure of the 2nd.. bu i think its (1/6)^5*(5/6).. Still you can confirm with someone..

OpenStudy (anonymous):

Wait, I misread the question.

OpenStudy (anonymous):

can you help me out with part (b)

OpenStudy (anonymous):

How many sides does the dice have? n?

OpenStudy (anonymous):

Or just 6?

OpenStudy (anonymous):

just 6

OpenStudy (anonymous):

So what do you mean by "the nth throw is n"?

OpenStudy (anonymous):

the number on the die is the same number as the throw

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

like first throw you get 1. i think thats what it means lol

OpenStudy (anonymous):

So for every throw there is a probability of 1/6 of getting the right number.

OpenStudy (akshay_budhkar):

Thomas9 i believe the first one is correct.. they are independent events.. everytime the probability is 1/6 so you just multiply the probabilities according to multiplication principle

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

first one is indeed correct.

OpenStudy (anonymous):

the first one is right but the second one isnt

OpenStudy (akshay_budhkar):

oh...

OpenStudy (anonymous):

i have the answer its 5/7776

OpenStudy (anonymous):

but im not sure how to get it

OpenStudy (anonymous):

It's binomial distributed, so you need to fill in the pdf. http://www.wolframalpha.com/input/?i=binomial+distribution+pdf

OpenStudy (anonymous):

n is the number of expirements, it's 6 in this case because we throw 6 times.

OpenStudy (anonymous):

p is the probability of succes, 1/6 in this case.

OpenStudy (anonymous):

and x is number of times of succes, in this case 5.

OpenStudy (anonymous):

\[P(X=5)=\left(\begin{matrix}6 \\ 5\end{matrix}\right)\frac{1}{6}^5\frac{5}{6}^{6-5}=6\frac{1}{6}^5\frac{5}{6}=\frac{5}{6^5}\]

OpenStudy (anonymous):

I hope that made sense to you.

OpenStudy (anonymous):

ok wait lol is there a multiplication inbetween each one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[P(X=5)=\left(\begin{matrix}6 \\ 5\end{matrix}\right)\cdot\frac{1}{6}^5\cdot\frac{5}{6}^{6-5}=6\cdot\frac{1}{6}^5\cdot\frac{5}{6}=\frac{5}{6^5}\]

OpenStudy (anonymous):

k thanks heaps:)